OpenStudy (anonymous):

Express in terms of i. √-16 + √-36

OpenStudy (shubhamsrg):

SQRT(-1) = i.. try now.. ain't that difficult

OpenStudy (callisto):

√-16 + √-36 = 4i + 6i = ..

OpenStudy (anonymous):

so then it would be 10i

OpenStudy (anonymous):

\[-16 = 4^2i^2\] \[-36 = 6^2i^2\] Put these values in square root brackets and then solve you will get your answer...

OpenStudy (anonymous):

then would it be 10^2i*i^2 ?

OpenStudy (anonymous):

ok i am lost sorry :(

OpenStudy (zzr0ck3r):

they are making this way to hard look at callisto

OpenStudy (anonymous):

6^2 = 36 and 4^2 = 16

OpenStudy (callisto):

To my knowledge \[\sqrt{-16} +\sqrt{ -36}\]\[=\sqrt{-1\times 4^2} +\sqrt{ -1 \times 6^2}\]\[=4\sqrt{-1} +6\sqrt{ -1}\]\[=4i +6i\]\[=10i\]

OpenStudy (zzr0ck3r):

yes, thats what a ti will tell you to:)

OpenStudy (anonymous):

then i did have it right 10i

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

thank you guys appreciate your help in understanding this kudos

OpenStudy (callisto):

welcome!